OpenStudy (anonymous):

what is the solution in this problem? The total annual income from capital invested at 5% is $380.40. The capital invested at 5.5% is $1,980 less than the capital invested at 5%. Find the amount of the two investments. representation: Let x= capital at 5.5% x + $1980 = capital at 5% (if reversed: capital at 5% is 1980 more than the capital at 5.5%) 0.055x = income from capital at 5.5% 0.05 (x+1980) = income from capital at 5%

OpenStudy (phi):

Did they give you the equations, or are they yours?

OpenStudy (anonymous):

its already given

OpenStudy (phi):

the last equation says 0.05 (x+1980) = income from capital at 5% they tell you in the first sentence income from capital invested at 5% is $380.40 so you can make that a real equation (with numbers on the right side) can you do that?

OpenStudy (anonymous):

can you please give me the real equation?

OpenStudy (phi):

the last equation says 0.05 (x+1980) = income from capital at 5% they tell you in the first sentence income from capital invested at 5% is $380.40

OpenStudy (anonymous):

so 0.05(x+1980)=380.40 yeah igot it thanks..

OpenStudy (phi):

now you need to solve it. Can you do that?

OpenStudy (anonymous):

yeah , thanks..